Give the precise thermochemical equation to which the following statements refer.
(i) The standard enthalpy of formation Of methane is - 74.8 K J/mol
(ii) The mean bond energy of C - H bond in methane is 413 KJ/mol
(iii) The enthalpy of atomization of carbon is 715 KJ/mol
1
Expert's answer
2020-12-03T13:48:11-0500

"1. \ C_{(s)}+2H_{2(g)} ⟶ CH_{4(g)}\quadΔH_f =−75kJ/mol"


"2." One C-H bond = 413kj/mol methane has 4 C-H bonds = 4(413) = 1652kj/mol


"CH_{4(g)} → C_{(g)} + 4H_{(g)} \quad ΔH° = 1652kJ/mol"


"3. \ C_{(s)} \to C_{(g)} \quad ΔH_{at} = 715kJ/mol"